Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 307     Entry time: 17 Aug 2023, 11:06
Author: Sergii Parchenko, Manuel Harder 
Group: SCS 
Subject: Night Shift Summary on 15.08.2023 

Achievements:

  1. Replaced focusing lens of PP laser and measured focus with knife edge, focus is about 720 micron vert. & 600 horizontal. Horizontal x-ray spot size is 220 micron
  2. Adapted fluence calibration for the new spot sizes
  3. Optimized spatial and temporal overlap with new lens
  4. Performed optical and x-ray intensity & sample damage tests. We found that up to 15% GATT and up to 50% laser power are okay to use, without altering the sample.
  5. We started to collect hRIXS data at positive (1ps) and negative (-2ps) delays.

Problems:

  1. Few beam dumps.
  2. BAM scene not working
  3. The new (?) centroiding function produces spectra with different intensity for acquisitions with the same settings. The old algorithm seems to work well.
ELOG V3.1.4-7c3fd00